OpenStudy (anonymous):

Dad is 4 times as old as his son Jim. In 10 years, Dad's age will be 20 years more than twice Jim's age. How old is Jim? answer list 10 years old 15 years old 30 years old

OpenStudy (anonymous):

15 years old

OpenStudy (anonymous):

@sauravshakya Prove it.

OpenStudy (anonymous):

how did u get that ya

OpenStudy (anonymous):

d=4j ....(i) 4j+10 = 2*(j+10)+20 .....(ii) solve these equations

OpenStudy (anonymous):

@sauravshakya Thank you for posting it. I wrote my equations wrong.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

your second equation is wrong after 10 years jim will be (j+10)years old

OpenStudy (radar):

By Jove, I overlooked that. 4j+10= 20+2(j+10) 4j+10=20 +2j+20 2j=30 j=15 now and dad is 60 now lets see if it checks out for ten years down the road j will be 25 and dad will 70 and sure enough dad is 20 plus twice junior

OpenStudy (radar):

Thanks sauravahakya for pointing that out, sorry johnjohn for the mathematical error (it won't be the last lol)